# Characteristic Root Analysis and Regularization for Linear Time Series Forecasting

### v1

@@ E.1 Dataset
We provide a summary of the dataset we used in Table 5. For ETT [Zhou et al., 2021] data, we follow previous works [Lin et al., 2024, Xu et al., 2023, Wu et al., 2022] and divide the dataset into train, val, and test sets with a $6:2:2$ ratio. For the rest of the datasets, they are divided in to train, validation, and test sets with a $7:1:2$ ratio.
@@
Additionally, we also use M4 [Makridakis et al., 2020] dataset to test the robustness of our methods under short lookback window as well as their performances in short-term forecasting tasks. The M4 dataset consists of 100,000 real-world time series from diverse domains such as finance, economics, demographics, and industry. Each subset of M4 is collected in yearly, quarterly, monthly, weekly, daily, and hourly frequencies, with a mixture of data from the domains mentioned before. For the M4 dataset, the lookback window $L$ is set to $2\times H$, where $H$ is the forecasting horizon, which differs from long-term forecasting datasets. The details of M4 dataset are listed in Tabel 6.
@@ E.2 Implementation Details
For more details on the running settings of Root Purge, please refer to our code and running scripts  [^9] Upon acceptance, we will open source our code..
@@ E.4 Comparing on Traffic and Electricity Dataset
For linear models, these large-scale datasets allow any linear model to reach the optimum via only the root-seeking loss. In Appendix E.8, we reaffirm that the optimal weight matrices, $\mathbf{W}$, are indeed of full rank. Therefore, it is expected that little or no regularization is needed given the dataset size and natural full-rankness of the time series.
As shown in Table 7, the performance of all linear models is very close on these two datasets. RRR and Root Purge either match or outperform other linear models (i.e., DLinear, FITS, and SparseTSF). Notably, for Root Purge, while the optimal setting is to use no regularizers and set $\lambda=0$, little performance degradation is observed with a relatively large regularization coefficient ($\lambda=0.125$). This reaffirms the robustness of our methods.
@@ E.6 Linear Forecaster with Bias and without Bias
ETTh1
96 $0.379\pm 0.001$ ${\color[rgb]{1,0,0}\mathbf{0.378}}\pm 0.001$ $0.375\pm 0.001$ ${\color[rgb]{1,0,0}\mathbf{0.374}}\pm 0.001$
@@
ETTh2
96 ${\color[rgb]{1,0,0}\mathbf{0.275}}\pm 0.001$ $0.276\pm 0.001$ ${\color[rgb]{1,0,0}\mathbf{0.272}}\pm 0.000$ $0.273\pm 0.000$
@@
ETTm1
96 $0.311\pm 0.002$ ${\color[rgb]{1,0,0}\mathbf{0.310}}\pm 0.001$ $0.310\pm 0.001$ ${\color[rgb]{1,0,0}\mathbf{0.308}}\pm 0.002$
@@
ETTm2
96 $0.163\pm 0.001$ ${\color[rgb]{1,0,0}\mathbf{0.163}}\pm 0.001$ $0.164\pm 0.000$ ${\color[rgb]{1,0,0}\mathbf{0.164}}\pm 0.000$
@@
Weather
96 $0.143\pm 0.000$ ${\color[rgb]{1,0,0}\mathbf{0.143}}\pm 0.001$ $0.145\pm 0.000$ ${\color[rgb]{1,0,0}\mathbf{0.145}}\pm 0.000$
@@
Exchange
96 $091\pm 0.001 0.0$${\color[rgb]{1,0,0}\mathbf{0.087}}\pm 0.0$$0$ $088\pm 0.001 0.0$${\color[rgb]{1,0,0}\mathbf{0.084}}\pm 0.0$$0$
@@
ECL
96 $0.134\pm 0.000$ ${\color[rgb]{1,0,0}\mathbf{0.134}}\pm 0.000$ $0.135\pm 0.000$ ${\color[rgb]{1,0,0}\mathbf{0.135}}\pm 0.001$
@@
Traffic
96 $0.386\pm 0.001$ ${\color[rgb]{1,0,0}\mathbf{0.386}}\pm 0.000$ $0.385\pm 0.000$ ${\color[rgb]{1,0,0}\mathbf{0.385}}\pm 0.000$
@@ E.7 Additional Empirical Results for Root Purge
Frequency Domain
ETTh1
96 0.374 0.373 0.371 0.368 0.364 0.359 0.354 0.351
@@
Time Domain
ETTh1
96 0.378 0.378 0.376 0.373 0.368 0.362 0.356 0.354
@@ E.8 Additional Empirical Results for Rank Reduction Methods
Figure 14: RRR Rank-MSE Trade-off on 4 example datasets. (a) & (b): On smaller datasets, there are clear benefits in truncating the rank of $\mathbf{W}$. (c) & (d): On a complex dataset with a large number of channels, there are many roots to be captured and $\mathbf{W}$ becomes full rank. In such a case, it is usually sufficient to use the root-seeking loss alone.
@@
In Figure 14, it can be seen that for smaller datasets such as ETTh1 and ETTh2, it is usually beneficial to truncate the rank, suggesting the existence of spurious roots as described in our theoretical analysis. Further, it is obvious to see that for Electricity and Traffic, it is optimal to choose full rank of $\mathbf{W}$, validating that for these large, complex datasets, it can be sufficient to use just the root-seeking loss alone. In addition, the full-rankness of $\mathbf{W}$ also suggests potential needs of higher model capacity, which might explain why non-linear models tend to perform better on these datasets.
@@
We leave the full plot with all datasets and all forecasting horizons $H=\{96,192,336,720\}$ at the end of the Appendix, in Figures 19, 21, 23, 25, 27, 29, 31, and $3$.
@@ E.10 Short-term Forecasting–An Extreme Case
Root Purge and RRR leverage rank and nullity manipulation to eliminate spurious roots. However, when a linear model is already underfitting and thus full-rank, no spurious roots are present. Consequently, we expect little to no performance gain from applying Root Purge or RRR to linear models on M4. On the other hand, our approaches—especially Root Purge—do not over-regularize the model even in these extreme cases (as shown in Table 17), suggesting their potential as a default regularizer for linear time series forecasting models.
@@ E.11 Higher-Order Root Purge
ETTh1
96 ${\color[rgb]{1,0,0}\mathbf{0.359}}\pm 0.000$ $0.365\pm 0.000$ $0.371\pm 0.000$ $0.373\pm 0.000$
@@
ETTh2
96 ${\color[rgb]{1,0,0}\mathbf{0.268}}\pm 0.000$ $0.270\pm 0.000$ $0.271\pm 0.000$ $0.271\pm 0.000$
@@
ETTm1
96 ${\color[rgb]{1,0,0}\mathbf{0.305}}\pm 0.000$ $0.306\pm 0.000$ $0.307\pm 0.000$ $0.308\pm 0.000$
@@
ETTm2
96 ${\color[rgb]{1,0,0}\mathbf{0.161}}\pm 0.000$ $0.161\pm 0.000$ $0.163\pm 0.000$ $0.163\pm 0.000$
@@
Weather
96 ${\color[rgb]{1,0,0}\mathbf{0.142}}\pm 0.000$ $0.143\pm 0.000$ $0.143\pm 0.000$ $0.144\pm 0.000$
@@
Exchange
96 ${\color[rgb]{1,0,0}\mathbf{0.082}}\pm 0.000$ $083\pm 0.000 0.0$$84\pm 0.000 0.0$$087\pm 0.002$
@@ Appendix G Declaration
Figure 31: Rank-MSE Trade-off Curves on Electricity
@@
Figure 33: Rank-MSE Trade-off Curves on Traffic

### v2

@@ D.7 Analysis of Root Purge
Remark 11.
For the stop-gradient objective
$\widetilde{J}(\mathbf{W})=\|\mathbf{Y}-\mathbf{XW}\|_{F}^{2}+\lambda\|\mathbf{R}\mathbf{W}\|_{F}^{2},$
where the residual matrix $\mathbf{R}$ is treated as constant during optimization, the gradient $\nabla\widetilde{J}(\mathbf{W})$ is Lipschitz continuous with constant $L$ satisfying:
$\widetilde{L}\leq 2\|\mathbf{X}\|_{2}^{2}+2\lambda(3M^{2}\|\mathbf{X}\|_{2}^{2}+4M\|\mathbf{X}\|_{2}\|\mathbf{Y}\|_{2}+\|\mathbf{Y}\|_{2}^{2})$ (12)
@@ E.1 Dataset
(a) ETTh1
(b) Weather
(c) Exchange Rate
(d) Traffic
Figure 12: Representative samples from the ETTh1, weather, exchange rate, and Traffic datasets, illustrating the complex, non-stationary trends characteristic of real-world time series data.
We provide a summary of the dataset we used in Table 6. For ETT [Zhou et al., 2021] data, we follow previous works [Lin et al., 2024, Xu et al., 2023, Wu et al., 2022] and divide the dataset into train, val, and test sets with a $6:2:2$ ratio. For the rest of the datasets, they are divided in to train, validation, and test sets with a $7:1:2$ ratio. The characteristics of these datasets are further illustrated by representative samples in Figure 12.
@@
Additionally, we also use M4 [Makridakis et al., 2020] dataset to test the robustness of our methods under short lookback window as well as their performances in short-term forecasting tasks. The M4 dataset consists of 100,000 real-world time series from diverse domains such as finance, economics, demographics, and industry. Each subset of M4 is collected in yearly, quarterly, monthly, weekly, daily, and hourly frequencies, with a mixture of data from the domains mentioned before. For the M4 dataset, the lookback window $L$ is set to $2\times H$, where $H$ is the forecasting horizon, which differs from long-term forecasting datasets. The details of M4 dataset are listed in Tabel 7.
@@ E.2 Implementation Details
For more details on the running settings of Root Purge, please refer to our code and running scripts in the supplementary material  [^9] Upon acceptance, we will open source our code..
@@ E.4 Comparing on Traffic and Electricity Dataset
For linear models, these large-scale datasets allow any linear model to approach the optimum via only the root-seeking loss. In Appendix E.8, we reaffirm that although the numerical analysis from RRR and DWRR still shows that the optimal rank of $\mathbf{W}$ is not full, the performance of full-rank models can come close to that of their rank-truncated counterparts.
As shown in Table 8, the performance of all linear models is very close on these two datasets. RRR and Root Purge either match or outperform other linear models (i.e., DLinear, FITS, and SparseTSF). Notably, for Root Purge, while data scaling has already pushed the model close to its optimality, performance remains robust with a relatively large regularization coefficient ($\lambda=0.125$). This reaffirms the robustness of our methods.
@@ E.6 Linear Forecaster with Bias and without Bias
ETTh1 96 $0.379\pm 0.001$ ${\color[rgb]{1,0,0}\mathbf{0.378}}\pm 0.001$ $0.375\pm 0.001$ ${\color[rgb]{1,0,0}\mathbf{0.374}}\pm 0.001$
@@
ETTh2 96 ${\color[rgb]{1,0,0}\mathbf{0.275}}\pm 0.001$ $0.276\pm 0.001$ ${\color[rgb]{1,0,0}\mathbf{0.272}}\pm 0.000$ $0.273\pm 0.000$
@@
ETTm1 96 $0.311\pm 0.002$ ${\color[rgb]{1,0,0}\mathbf{0.310}}\pm 0.001$ $0.310\pm 0.001$ ${\color[rgb]{1,0,0}\mathbf{0.308}}\pm 0.002$
@@
ETTm2 96 $0.163\pm 0.001$ ${\color[rgb]{1,0,0}\mathbf{0.163}}\pm 0.001$ $0.164\pm 0.000$ ${\color[rgb]{1,0,0}\mathbf{0.164}}\pm 0.000$
@@
Weather 96 $0.143\pm 0.000$ ${\color[rgb]{1,0,0}\mathbf{0.143}}\pm 0.001$ $0.145\pm 0.000$ ${\color[rgb]{1,0,0}\mathbf{0.145}}\pm 0.000$
@@
Exchange 96 $091\pm 0.001 0.0$${\color[rgb]{1,0,0}\mathbf{0.087}}\pm 0.0$$0$ $088\pm 0.001 0.0$${\color[rgb]{1,0,0}\mathbf{0.084}}\pm 0.0$$0$
@@
ECL 96 $0.134\pm 0.000$ ${\color[rgb]{1,0,0}\mathbf{0.134}}\pm 0.000$ $0.135\pm 0.000$ ${\color[rgb]{1,0,0}\mathbf{0.135}}\pm 0.001$
@@
Traffic 96 $0.386\pm 0.001$ ${\color[rgb]{1,0,0}\mathbf{0.386}}\pm 0.000$ $0.385\pm 0.000$ ${\color[rgb]{1,0,0}\mathbf{0.385}}\pm 0.000$
@@ E.7 Additional Empirical Results for Root Purge
Frequency Domain ETTh1 96 0.374 0.373 0.371 0.368 0.364 0.359 0.354 0.351
@@
Time Domain ETTh1 96 0.378 0.378 0.376 0.373 0.368 0.362 0.356 0.354
@@ E.8 Additional Empirical Results for Rank Reduction Methods
Figure 15: RRR Rank-MSE Trade-off on 4 example datasets. (a) & (b): On smaller datasets, there are clear benefits in truncating the rank of $\mathbf{W}$. (c) & (d): On larger datasets (Electricity, Traffic), the performance curve is robust to the rank of $\mathbf{W}$, and a full-rank matrix does not cause significant performance degradation—a result of data scaling, where sufficient data helps suppress spurious roots. However, although the curves for these larger datasets appear flat at this scale, numerical analysis reveals that the optimal model is not achieved at full rank, indicating that slight rank truncation remains beneficial despite data scaling.
@@
In Figure 15, it can be seen that for smaller datasets such as ETTh1 and ETTh2, it is usually beneficial to truncate the rank, suggesting the existence of spurious roots as described in our theoretical analysis. Conversely, for the larger and more complex Electricity and Traffic datasets, a full-rank $\mathbf{W}$ does not significantly degrade performance, which aligns with the data scaling effect in Proposition 1: sufficient data suppresses spurious roots. That said, our experiments with RRR and DWRR reveal that even for the Electricity and Traffic datasets, the optimal model is not at full rank; slight truncation remains beneficial despite data scaling. This indicates that a low-rank structure is relatively common in time series forecasting. Consequently, applying regularization by default can be a generally beneficial practice for time series forecasting tasks.
@@
We leave the full plot with all datasets and all forecasting horizons $H=\{96,192,336,720\}$ at the end of the Appendix, in Figures 20, 22, 24, 26, 28, 30, 32, and 34.
@@ E.10 Short-term Forecasting–An Extreme Case
Root Purge and RRR leverage rank and nullity manipulation to eliminate spurious roots. However, when a linear model is already significantly underfitting, no spurious roots are present. Consequently, we expect little to no performance gain from applying Root Purge or RRR to linear models on M4. On the other hand, our approaches—especially Root Purge—do not over-regularize the model even in these extreme cases (as shown in Table 18), suggesting their potential as a default regularizer for linear time series forecasting models.
@@ E.11 Higher-Order Root Purge
ETTh1 96 ${\color[rgb]{1,0,0}\mathbf{0.359}}\pm 0.000$ $0.365\pm 0.000$ $0.371\pm 0.000$ $0.373\pm 0.000$
@@
ETTh2 96 ${\color[rgb]{1,0,0}\mathbf{0.268}}\pm 0.000$ $0.270\pm 0.000$ $0.271\pm 0.000$ $0.271\pm 0.000$
@@
ETTm1 96 ${\color[rgb]{1,0,0}\mathbf{0.305}}\pm 0.000$ $0.306\pm 0.000$ $0.307\pm 0.000$ $0.308\pm 0.000$
@@
ETTm2 96 ${\color[rgb]{1,0,0}\mathbf{0.161}}\pm 0.000$ $0.161\pm 0.000$ $0.163\pm 0.000$ $0.163\pm 0.000$
@@
Weather 96 ${\color[rgb]{1,0,0}\mathbf{0.142}}\pm 0.000$ $0.143\pm 0.000$ $0.143\pm 0.000$ $0.144\pm 0.000$
@@
Exchange 96 ${\color[rgb]{1,0,0}\mathbf{0.082}}\pm 0.000$ $083\pm 0.000 0.0$$84\pm 0.000 0.0$$087\pm 0.002$
@@ Appendix G Declaration
Figure 32: Rank-MSE Trade-off Curves on Electricity. Although the curves for these larger datasets appear flat at this scale, numerical analysis reveals that the optimal model is not achieved at full rank, indicating that slight rank truncation remains beneficial despite data scaling.
@@
Figure 34: Rank-MSE Trade-off Curves on Traffic. Although the curves for these larger datasets appear flat at this scale, numerical analysis reveals that the optimal model is not achieved at full rank, indicating that slight rank truncation remains beneficial despite data scaling.